# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ 2.2 Our Mixed-Methods Approach
Our mixed-methods analytical framework offers a generalizable approach for studying human discretion and expertise in other settings involving algorithmically assisted decision making. The combination of causal inference, statistical testing, and qualitative analysis provides a holistic set of tools that extend beyond college advising to domains such as healthcare, social services, and hiring, where understanding the interaction between human judgment and algorithmic support is equally critical.
@@
As researchers, our identities, disciplinary training, and institutional affiliations shape how we approached the research questions, interpreted the data, and interacted with the broader context of advising and algorithmic decision making. The core research team includes a graduate student and two tenure-track faculty at R1 institutions (Authors 1, 4, and 5), with training in computer science, operations research, and the societal impacts of algorithmic decision making in educational applications. The research process was shaped by a collaboration between the core research team and staff at the National Institute for Student Success (NISS) at GSU. From November 2023 to November 2024, Authors 4 and 5 held biweekly meetings with staff members at NISS, including Authors 2 and 3, to understand the MAAPS intervention and its implementation context, as well as to determine the structure and limitations of the advising data. These meetings grounded our quantitative and qualitative analyses in the realities of the program.
@@
Our access to the MAAPS data is limited to de-identified records shared under a data use agreement with NISS, and we have not interacted directly with students or advisors. The Princeton University and Columbia University Institutional Review Boards determined the research was exempt from further review, as it does not qualify as human subjects research. Our interpretation of advising practices is shaped by our outsider perspective and by the structure and limitations of the available data. In the qualitative analysis, we sought to preserve the nuance of advisors’ written reflections while acknowledging the constraints of interpreting them at a distance. Finally, because we conduct a secondary analysis of the MAAPS experiment, the original design and data availability have also shaped our research questions and methods.
3 Theoretical Framework: Expertly Targeted Interventions
@@
A student meets with their advisor, explaining that a medical condition prevents them from walking long distances. In response, the advisor works with the student to design a schedule with nearby classrooms that also meets their academic interests. The student is able to achieve a high semester GPA.
This example, based on a real case, illustrates how advisors can draw on rich, algorithmically “invisible” information to support students in achieving their goals—in this case, a health condition that impacts the student’s schedule. These insights create an information asymmetry between human decision makers and algorithmic tools, raising the question: what role does human discretion play in delivering interventions and improving outcomes?
In this section, we propose a causal framework to understand the role of human expertise in improving outcomes when algorithmic decision aides are deployed in interventional settings, such as the MAAPS collegiate advising case.
We begin in §3.1 by introducing our model of algorithm-assisted intervention, which encapsulates the model of Alur et al. [4] for prediction as a special case. §3.2 defines the concept of human expertise in targeted actions and specifies assumptions about the causal mechanism. Finally, §3.3 provides theoretical results on detecting the presence expertly targeted actions from data. We show that, without further assumptions, it is impossible to detect expertly targeted actions from observational quantitative data alone. However, we define statistical conditions that can provide important evidence suggestive of expertly targeted action.
3.1 Model of Expertise in Interventions
@@
We begin by providing a causal model of algorithm-assisted decision-making processes, and a causal definition of expertise in this setting. We use the machinery of structural causal models (SCMs), which define the way that variables causally influence one another through a set of functional relationships [49, 50, 53, 9].
We consider a model where a human decision-maker, at each time step $t$, needs to make a decision—corresponding in the MAAPS study to the action taken by an advisor in student meetings.
Let $\mathbf{X}_{t}$ be the algorithmically-observed features measured at time $t$. In the MAAPS experiment, $\mathbf{X}_{t}$ refers to the observable administrative data on the student. Let $Y_{t+1}$ denote the outcome variable of interest at the next time step, for instance, the student’s term GPA. We define $\mathbf{U}_{t}$ as the context at time $t$ that is relevant to the outcome $Y_{t+1}$ but unobserved by the algorithm; for instance the information learned through advising conversations. Here, outcome relevance means that $\mathbf{U}_{t}$ is an input of the process generating the student’s outcomes $Y_{t+1}$, to be formalized shortly. [^6] If there is no such relevant context, then $\mathbf{U}_{t}$ may be excluded from $\mathbf{V}_{\mathcal{A}}$.
The advisor observes $\mathbf{X}_{t}$ and potentially $\mathbf{U}_{t}$ and then takes action $A_{t}$ (e.g., advising the student to talk to their instructor).
[…]
For each variable $V\in\mathbf{V}_{\mathcal{A}}$, we define a function $f_{V}\in\mathcal{F}_{\mathcal{A}}$ that models $V$ as the causal output of a subset of other variables in $\mathbf{V}_{\mathcal{A}}$, together with an exogenous noise variable $\epsilon_{V}\in\mathbf{\epsilon}_{\mathcal{A}}$ specific to $V$.
We assume that these functions satisfy the input–output constraints specified by the directed acyclic graph (DAG) in Figure 1, where a directed edge from variable $V_{1}$ to $V_{2}$ indicates that $f_{V_{2}}$ takes $V_{1}$ as an input. For example, the structural equation $Y_{t+1}=f_{Y}(\mathbf{X}_{t},\mathbf{U}_{t},A_{t})$ corresponds to a causal graph in which $\mathbf{X}_{t}$, $\mathbf{U}_{t}$, and $A_{t}$ are direct causes (parents) of $Y_{t+1}$. [^8] Note that when no endogenous context $\mathbf{U}_{t}$ is in the model, i.e. $\mathbf{U}_{t}\notin\mathbf{V}_{\mathcal{A}}$, then $\mathbf{U}_{t}$ will not be a causal parent of $Y_{t+1}$.
Our model generalizes the functional model of human expertise in prediction, presented in Alur et al. [4], in two ways. First, we fully specify the causal structure among variables using DAGs. Second, we move beyond prediction to intervention, accounting for the possibility that the human decision-maker’s choices can influence future outcomes. [^9] Specifically, we replace the prediction variable $\hat{Y}$ [see eq. 5, 4]—which was assumed to have no effect on the predictive target $Y$—by an action variable $A_{t}$. The action $A_{t}$ may have a causal effect on the future outcome $Y_{t+1}$. Also, the action $A_{t}$ may be informed by both an algorithmic signal (represented by $\mathbf{X}_{t}\to A_{t}$) and additional, unobserved contextual knowledge (represented by $\mathbf{U}_{t}\to A_{t}$). [^10] Our model can capture three types of algorithm-assisted decision making of increasing generality. In the special case $A_{t}$ is performing an idle prediction $\hat{Y}$, we can simply stipulate $A_{t}$ does not influence the outcome $Y_{t+1}$ and our model reduces to Alur et al. [4]’s model. But unlike Alur et al. 4, our model can also represent predictive decision making, where predictions are made of an outcome in order to influence that outcome. With $R(X)$ denoting the range of a random variable $X$, we may think of $A$ as determined by a (potentially noisy) decision policy $\pi:R(\hat{Y})\to R(A)$ on the human’s internal predictions $\hat{Y}$. Finally, our model can represent more general algorithm-assisted interventional contexts, in which actions to affect an outcome do not necessarily depend on predictions of that outcome. Collegiate advising is one such example: An advisor may decide to take actions to support students for many reasons that are not all well-modeled as predictions of the outcome they aim to influence.
[…]
@@
Finally, the model specifies a joint distribution $\mathbb{P}$ over the exogenous noise variables $\mathbf{\epsilon_{\mathcal{A}}}$, assumed to be independent. Together with the structural equations, this distribution induces unique observational and causal distributions on all model variables. The observational distribution is Markovian with respect to the causal DAG. [^11] That is, every variable is independent of its non-descendants given its parents. As is standard for causal identification from observational data, we assume the model is causally minimal, that is, the distribution is not Markovian with respect to any proper subgraph [53, pp. 107-109].
@@ 3.2 Definition of Expertly Targeted Action
Say a MAAPS advisor chooses to recommend a particular course to a student in a meeting. Given a full causal model of the decision-making scenario, when does this course recommendation demonstrate expertise? Here are three intuitive requirements:
@@
    The recommendation helps the student’s outcomes, e.g. their GPA. For instance, it should not be that the student ends up performing worse than they would have without taking the course.
@@
    The recommendation was targeted based on outcome-relevant information unavailable to the decision support algorithm —such as insights from a conversation about the student’s interests. For example, it should not be that the advisor recommends this course to every student, or that the recommendation is determined by the algorithm’s output.
@@
    The recommendation is effective because of the targeting. For instance, it should not be that all students who take this course get an A+ (and so the targeting was unimportant to the effectiveness of the intervention), but that the course recommendation uniquely benefits the targeted student.
These requirements can be stated as three causal statements—one about the action’s effect, one about the action’s targeting, and one about their interaction. When the three are true together, we call that expertly targeted action. To simplify the statement of the definition, we consider a setting in which both the action $A_{t}$ and external information $U_{t}$ are binary variables. [^12] We discuss in Appendix A.2.3 how to generalize the definition to accommodate non-binary contrasts.
Definition 3.1 (Expertise).
For a given observation with $\mathbf{X}_{t}=\mathbf{x}$, $U_{t}=u$, and $A_{t}=1$, a human decisionmaker’s action is expertly targeted when the following three conditions are met:
  1. 1.
    Effective action. Taking the action improves the outcome.
    $\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)>\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
  2. 2.
    Targeted action. The human decisionmaker chose to take the action based on non-algorithmic information.
    $\mathbb{P}(A_{t}=1|do(U_{t}=u),\mathbf{X}_{t}=\mathbf{x})>\mathbb{P}(A_{t}=1|do(U_{t}=1-u),\mathbf{X}_{t}=\mathbf{x})$
  3. 3.
@@
    Heterogeneous action. The relative effectiveness of the action is due to heterogeneity in $U$.
    $\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=1),\mathbf{X}_{t}=\mathbf{x})$ $-\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=0),\mathbf{X}_{t}=\mathbf{x})>$
    $\mathbb{P}(Y_{t+1}=1|do(U_{t}=1-u,A_{t}=1),\mathbf{X}_{t}=\mathbf{x})$ $-\mathbb{P}(Y_{t+1}=1|do(U_{t}=1-u,A_{t}=0),\mathbf{X}_{t}=\mathbf{x})$
To capture instances of expertise according to this definition, it is important to describe actions at the right level of abstraction. For instance, say the action, $A_{t}$, is defined as “recommending a schedule.” If advisors recommends all students a schedule, then the targeted action component of the expertise definition will fail—even if the advisors are incorporating non-algorithmic information into the particular schedules they recommend to students.
3.3 Impossibility of Identifying Expertise from (Quantitative) Observational Data
@@
The above definition considers how we would identify instances of expertise with full access to the causal mechanisms. However, in most cases we will only have an observational dataset $\mathcal{D}$ of $n$ $(\mathbf{x_{i}},y_{i},\hat{y}_{i})$, observations drawn from the joint distribution over variables $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$. [^13] This observational distribution excludes $U_{t}$ because it is uncommon to collect data on $U_{t}$, the private, algorithmically-invisible information that humans incorporate in their decision making (and moreover, it is unclear how to systematically collect it). If we do not make any causal assumptions beyond those in §3.1 and Figure 1, can we identify whether an advisor demonstrates expertise in targeting from this observational information?
@@
We prove an impossibility result that we cannot detect human expertise from observational data alone, because causal models that differ on the relevant causal facts can be statistically indistinguishable. However, we show that a related statistical condition—conditional dependence of actions $A_{t}$ and future outcomes $Y_{t+1}$ given algorithmically available data $\mathbf{X}_{t}$—is a necessary condition for expertly targeted action, and a sufficient condition for weakened components of it. We then test whether this condition is met in the MAAPS dataset in §4.

### v3

@@ 3.2 Our Mixed-Methods Approach
Our domain-agnostic approach, combining causal inference, statistical testing, and qualitative analysis, applies to studying human discretion and expertise across domains, from college advising to healthcare, social services, hiring, and other settings where understanding the interaction between human judgment and algorithmic support is critical.
@@
As researchers, our identities, disciplinary training, and institutional affiliations shape how we approached the research questions, interpreted the data, and interacted with the broader context of advising and algorithmic decision making. The core research team (Authors 1, 4, and 5) included a graduate student and two tenure-track faculty at R1 institutions, with training in computer science, operations research, and the societal impacts of algorithmic decision making in educational applications in various U.S.-based contexts. The research process was shaped by a collaboration between the core research team and staff at the National Institute for Student Success (NISS) at GSU. From November 2023 to November 2024, Authors 4 and 5 held biweekly meetings with staff members at NISS, including advising Authors 2 and 3, to understand the MAAPS intervention and its implementation context, as well as to determine the structure and limitations of the advising data. These meetings grounded the quantitative and qualitative analyses in the realities of the program.
@@
Access to the MAAPS data was limited to de-identified records shared under a data use agreement with NISS, and the core research team did not interact directly with students or advisors. The Princeton University and Columbia University Institutional Review Boards determined the research was exempt from further review, as it does not qualify as human subjects research. The interpretation of advising practices was shaped by the core team’s outsider perspective despite best efforts to understand the unique institutional context at GSU, and by the structure and limitations of the available data. In the qualitative analysis, the team sought to preserve the nuance of advisors’ written reflections while acknowledging the constraints of interpreting them at a distance.
Finally, as the current article constitutes a secondary analysis of the MAAPS experiment, the original design and data availability have also shaped current research questions and methods.
4 Theoretical Framework: Expert Interventions
@@
Consider the following intuitive example ofalgorithmically-“invisible” information guiding discretionary action in advising: A student meets with their advisor, explaining that a medical condition prevents them from walking long distances. In response, the advisor works with the student to design a schedule with nearby classrooms that also meets their academic interests. The student is able to achieve a high semester GPA.
Based on a real case in the MAAPS study, this example illustrates how advisors can draw on rich, “non-algorithmic” context to target their support—in this case, a health condition that impacts the student’s schedule. Crucially, human expertise in the interventional setting involves two components: (1) accessing this additional context and (2) also appropriately acting on it through tailored guidance to positively influence student outcomes. This raises the question: what role does this uniquely human expertise—using contextual knowledge to guide appropriate action—play in improving outcomes?
In this section, we propose a causal framework to answer that question. We present a stylized model of human actions in the presence of algorithmic decision aids, and resulting outcomes. The model is inspired by and generalizes Alur et al.’s [6] framework for detecting human expertise in the purely predictive setting. While their framework address the first requirement of (1) accessing the additional context to better predict the student’s state, we extend this to include the second requirement of (2) acting appropriately on the context to improve outcomes through interventions. [^6] Alur et al. [6] make the assumption that predictions are not “performative” [67], that is, that they do not impact outcomes. We begin in §4.1 by introducing our model for algorithm-assisted interventions. §4.2 defines human expertise in action targeting as the use of non-algorithmic context to select more effective actions. In §4.3, presents theoretical results on detecting the presence of expertly targeted actions from data. Finally, §4.4 discusses how this theoretical framework informs our quantitative and qualitative investigation of the MAAPS data.
4.1 Model of Decision Making
@@
We model algorithm-assisted decision-making processes using structural causal models (SCMs), which define the way that variables causally influence one another through a set of functional relationships [64, 65, 69, 13].
Consider a scenario where a human decision maker, at each time step $t$, needs to make a decision to try to improve outcomes at time $t+1$. We define an SCM $\mathcal{A}=\langle\mathbf{V}_{\mathcal{A}},\mathbf{\epsilon}_{\mathcal{A}},\mathcal{F}_{\mathcal{A}},\mathbb{P}_{\mathcal{A}}\rangle$ with endogenous variables $\mathbf{V}_{\mathcal{A}}$, exogenous noise $\mathbf{\epsilon}_{\mathcal{A}},$ structural equations $\mathcal{F}_{\mathcal{A}}$ and probability distribution $\mathbb{P}_{\mathcal{A}}$ to represent this single-stage decision scenario. $\mathbf{V}_{\mathcal{A}}$ includes $A_{t}$, the action taken by the advisor, and $Y_{t+1}$, the outcome at the next time step, like the student’s term GPA, and may also include $\mathbf{X}_{t}$, any outcome-relevant administrative data about students observed by advisors (through an algorithmic decision aid), and $\mathbf{U}_{t}$, any outcome-relevant non-algorithmic context at time $t$ (for instance, the information learned through advising conversations). [^7] If there is no algorithmically-available data observed and incorporated by advisors, then $\mathbf{X}_{t}$ may be excluded from $\mathbf{V}_{\mathcal{A}}$. If there is no non-algorithmic context observed and incorporated by advisors, then $\mathbf{U}_{t}$ may be excluded from $\mathbf{V}_{\mathcal{A}}$. In these cases, Assumption 4.1 is adjusted to omit assumptions about edges between excluded variables and any other model variables.
Here, outcome relevance means that a variable is an input of the process generating the student’s outcomes $Y_{t+1}$. These processes are defined by functions $f_{V}\in\mathcal{F}_{\mathcal{A}}$ for each variable $V\in\mathbf{V}_{\mathcal{A}}$, modeling $V$ as the causal output of a subset of other variables in $\mathbf{V}_{\mathcal{A}}$, and exogenous noise $\epsilon_{V}\in\mathbf{\epsilon}_{\mathcal{A}}$.
The causal model $\mathcal{A}$ induces a directed acyclic graph (DAG) $\mathcal{G}_{\mathcal{A}}$ representing these processes, with nodes corresponding to endogenous model variables $\mathbf{V}_{\mathcal{A}}$, and a directed edge from variable $V_{1}$ to $V_{2}$ whenever $f_{V_{2}}$ takes $V_{1}$ as an input. We make an assumption about the functional relationships $\mathcal{F}_{\mathcal{A}}$ by imposing constraints on $\mathcal{A}$’s induced graph:
Assumption 4.1 (Graphical constraints).
As depicted in Figure 1, $\mathcal{G}_{\mathcal{A}}$ includes directed edges from $\mathbf{X}_{t}$ and $\mathbf{U}_{t}$ to $Y_{t+1}$, may include directed edges from $\mathbf{X}_{t}$ and $\mathbf{U}_{t}$ to $A_{t}$, may include a directed edge from $A_{t}$ to $Y_{t+1}$, and has no other directed edges.
This means the outcome is a process of both the algorithmically-available features ($\mathbf{X}_{t}\to Y_{t+1}$) and non-algorithmic context ($\mathbf{U}_{t}\to Y_{t+1}$). Advisors take actions that may be based on both these sources of information ($\mathbf{X}_{t}\to A_{t}$ and $\mathbf{U}_{t}\to A_{t}$). These actions may then influence the student’s future outcomes ($A_{t}\to Y_{t+1}$).
@@
Finally, the model specifies a joint distribution $\mathbb{P}_{\mathcal{A}}$ over the exogenous noise variables $\mathbf{\epsilon_{\mathcal{A}}}$. Together with the structural equations $\mathcal{F}_{\mathcal{A}}$, $\mathbb{P}_{\mathcal{A}}$ induces unique observational and causal distributions on all model variables.
@@ 4.2 Definition of Expertly Targeted Action
Suppose a MAAPS advisor recommends a particular course to a student. When does this action demonstrate expertise? We state three criteria that must be met for the action to count as expertly chosen. To simplify the definition, we consider a setting in which both the action $A_{t}$ and external information $U_{t}$ are binary variables, discussing generalizations to nonbinary contrasts in Appendix A.4.
Definition 4.1 (Expertly targeted action).
For a student with $\mathbf{X}_{t}=\mathbf{x}$, $U_{t}=u$, a human decision maker’s action $A_{t}=1$ is expertly targeted when three conditions are met:
@@
    Effective action. Taking the action improves the outcome. (Requires $A_{t}\to Y_{t+1}$).
    $\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=1),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$ $>$
    $\mathbb{P}(Y_{t+1}=1\mid do(A_{t}=0),\mathbf{X}_{t}=\mathbf{x},U_{t}=u)$
@@
    Targeted action. The human decisionmaker chose to take the action based on non-algorithmic information. (Requires $U_{t}\to A_{t}$).
    $\mathbb{P}(A_{t}=1|do(U_{t}=u),\mathbf{X}_{t}=\mathbf{x})$ $>$
    $\mathbb{P}(A_{t}=1|do(U_{t}=1-u),\mathbf{X}_{t}=\mathbf{x})$
@@
    Heterogeneous effects. The relative effectiveness of the action is driven by heterogeneity in the non-algorithmic information.(Requires both $A_{t}\to Y_{t+1}$ and $U_{t}\to A_{t}$).
    With $\Delta(u)=\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=1),\mathbf{X}_{t}=\mathbf{x})-\mathbb{P}(Y_{t+1}=1|do(U_{t}=u,A_{t}=0),\mathbf{X}_{t}=\mathbf{x})$,
    $\Delta(u)$ $>\Delta(1-u)$
To count as expertly targeted, then, a course recommendation should improve the student’s GPA (effective action). The advisor should choose to recommend the course based on context unavailable to an algorithm, say the student’s interests and health needs as learned in an advising conversation (targeted action). And the course should improve the student’s GPA because it meets the student’s interests, not because, say, all students in the course get As (heterogeneous effects). [^8] This definition of expertise reflects a notion of targeting as personalizing interventions based on causal heterogeneity [8, 31, 83]. Similar to a definition of “actionable heterogeneity” proposed by Shchetkina and Berman [83], we require that heterogeneity in the targeted attribute contributes to observed improvements from the action (heterogeneous effects). However, unlike Shectkina and Berman’s, our framework does not require that the ideally rational decision maker should not have performed the action if the individual did not possess the targeted attribute; we replace this with a requirement that the actual decision maker would not have performed the action if the individual did not have this attribute (targeted action). This choice corresponds to a difference in setting. Our goal is to identify instances of targeting in an individual agent’s behavior, rather than identify whether ideal policies can benefit from targeting, as Shchetkina and Berman do.
4.3 Identifying Expertly Targeted Action
@@
Suppose we have an observational dataset $\mathcal{D}$ of $n$ $(\mathbf{x_{i}},y_{i},\hat{y}_{i})$, observations drawn from the joint distribution over variables $\mathbf{X}_{t}$, $A_{t}$ and $Y_{t+1}$. [^9] This observational distribution excludes $U_{t}$ because it is uncommon to collect data on $U_{t}$, the private, algorithmically-invisible information that humans incorporate in their decision making (and moreover, it is unclear how to systematically collect it). Under the definitions in §4.1, can we identify whether an advisor demonstrates expertise in targeting from this observational data?
@@
We prove an impossibility result that we cannot detect human expertise from observational data alone. However, a related statistical condition—conditional dependence of actions $A_{t}$ and future outcomes $Y_{t+1}$ given algorithmically available data $\mathbf{X}_{t}$—is a necessary condition for expertly targeted action, and sufficient for related criteria. We test whether this condition is met in the MAAPS dataset in §5.
4.3.1 Assumptions
In our theoretical results, we invoke three common assumptions about $\mathbb{P}_{\mathcal{A}}$, the distribution over the noise variables $\mathbb{\epsilon_{\mathcal{A}}}$:
Assumption 4.2 (Markovian condition).
The distribution $\mathbb{P}_{\mathcal{A}}$ is such that the noise variables $\mathbb{\epsilon_{\mathcal{A}}}$ are jointly independent [65, p.30].
Assumption 4.3 (Causal minimality).
The distribution $\mathbb{P}_{\mathcal{A}}$ is not Markovian with respect to any proper subgraph of $\mathbf{V}_{\mathcal{A}}$.
Assumption 4.4 (Faithfulness).
For all disjoint subsets of variables $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}_{\mathcal{A}}$, $\mathbf{X}\perp\mathbf{Y}|\mathbf{Z}\implies\mathbf{X}\perp_{\mathcal{G}_{\mathcal{A}}}\mathbf{Y}\ |\ \mathbf{Z}$, where $\perp_{\mathcal{G}_{\mathcal{A}}}$ is the relation of d-separation relative to $\mathcal{G}_{\mathcal{A}}$.
The Markovian condition is a standard assumption that rules out hidden confounding between model variables [13, p.35]. Assuming the SCM is Markovian ensures the SCM’s DAG meets the global Markov condition for graphs, that d-separation implies conditional independence [69, pp.101-102, p.105]. Though standard, the absence of hidden confounding is a relatively strong assumption. [^10] We consider potential violations of this Markovian condition in our application of the theoretical framework to quantitative and qualitative analysis of expertise in the MAAPS experiment. Causal minimality, by contrast, is a weak assumption, sometimes viewed as a convention, that rules out models with redundant descriptions of causal relationships [69, pp. 107-109].
Faithfulness, our strongest assumption, is the converse of the global Markov condition [69, p.107] and implies causal minimality [69, p.108]. Unlike causal minimality, it conveys a substantive assumption about the causal mechanisms, that causal effects do not perfectly “cancel” each other out so as to make causally related variables appear independent [69, pp.107-108]. For instance, in our model, it could be violated if $A_{t}\leftarrow\mathbf{U}_{t}\to Y_{t+1}$ perfectly offset the effect $A_{t}\to Y_{t+1}$. Perfect cancellation is rare, justifying the assumption [69, p.136]. [^11] For instance, for linear models, such “cancelling” happens with zero probability when coefficients are drawn randomly [69, 86]. Since it allows for identifying graph structure from statistical independence, it is also a common assumption for causal inference from observational data, though not as common as assuming causal mechanisms are Markovian. We note only one of our results (Proposition 4.2) depends on it.